1. Consider a family where both the husband and wife work, and assume each spouse has three more years to work (after the current year) before retirement. The table below indicates the projected combined salary over the remaining work years if they each stay with their current employers. It also shows the combined salary they can expect if they relocate and change employers.

Year

Combined Salary at Current Jobs

Combined Salary at New Jobs

1

$80,000

$83,000

2

$82,000

$87,000

3

$85,000

$92,000

Suppose the total cost of moving, including direct expenses, foregone earnings, and psychic costs total $10,000. Assume all costs are incurred during the current year.

If the discount rate is equal to the market interest rate of 6%, is this investment in mobility worthwhile?

What is the maximum mobility costs could be and still make this investment worthwhile?

How reasonable is it to assume that the psychic costs of mobility all occur in the current year?

Is it reasonable to expect that there will also be psychic benefits associated with moving? If so, give some examples.

When a person moves without a new job already in hand, evidence indicates that the mobility may initially lead to lower earnings than in the current job. This reduction is then more than overcome in future periods by faster earnings growth in the new location. Holding all else constant, what is the lowest the new combined salary could be in year 1 and still have this be a worthwhile investment?
Answer #1

A) is this investment in mobility worthwhile?

Answer:-

the total cost of moving, including direct expenses, foregone earnings, and psychic costs total = $10,000

at T=0 that Means for higher earnings intial investement required = $ 10,000 ( cash Outflow)

At t = Additional Income

1 83000-80000 = $ 3000

2 87000-82000 = $5000

3 92000-85000 = $7000

Discounting These Cash flows @ 6% at t= 0

1. $ 3000 / (1+0.06) = $2830.20

2. $5000/(1+0.06)2 = $4449.98

3. $ 7000/ (1+0.06)3 = $ 5877.34

Total Cash Inflows at t=0 = $ 13,157.52

Therefore Net Present Value will be Inflows - outflows

= $13,157.52- 10,000 = $ 3,157.52

Mobolity is Viable from NPV technique Prespective.

b) What is the maximum mobility costs could be and still make this investment worthwhile?

Answer:-

To make the mobility possible one should get much amount of earning which they already earns

that means if cash inflows = cash outflows this will make no change in their regular earnings

so maximum mobility cost = $ 13,157.52

C) How reasonable is it to assume that the psychic costs of mobility all occur in the current year?

Answer:-

From my prespective psychic costs represents costs incurred being getting familier to new work space,

new living environment and other related costs.

when a person switch to new job area or new environment they have to face psychic problems so it is reasonable to assume that psychic costs will incurred in the current year because it depends when the person moves to new area , after some times he will become familier and get adapted to new environment.

d) Is it reasonable to expect that there will also be psychic benefits associated with moving? If so, give some examples?

Answer:-

It might be possible that moving to new environment might give some psychic benefits to a person.

such as :- Good Locality, Friendly Neighbours, creative workspace enviornment and enhanced job satisfaction.